# Sample-Specific Noise Injection For Diffusion-Based Adversarial Purification

### v1

@@ Appendix I Ablation Study on Bias Term
[…]
@@
[…]
@@
[…]
@@
Suppose there exists a constant $K>0$ such that for all $t\geq 0$ and all ${\mathbf{x}}_{t}\in{\mathcal{X}}$, the expected norm of the clean data given ${\mathbf{x}}_{t}$ satisfies: $\mathbb{E}_{{\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0}|{\mathbf{x}})}[\left\|{\mathbf{x}}_{0}\right\|]\leq K\left\|{\mathbf{x}}_{t}\right\|$. Then, there exist constants $10<C<$ 1 and $T_{0}>0$ such that for all $t\geq T_{0}$:
@@
We then turn to look into the asymptotic behavior of $g(t)$. We have $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. Since $10<1-\beta_{s}<$ 1 for all $s$, and the sequence is decreasing (as $\beta_{s}$ is increasing), it is easy to check that
@@
11 11 $>C.$
Putting them together, we have $\epsilon\in(0,\epsilon_{1})\cup(\epsilon_{2},\infty)$ that makes $\left\|s_{t}({\mathbf{x}})\right\|\geq C\left\|{\mathbf{x}}\right\|$ hold, given a constant $10<C<$ 1 in relation to $K$ and $\epsilon$. Setting $T_{0}=T_{\epsilon}$ completes the proof. ∎
@@
[…]
@@
Let $C_{1}=\sup_{t\geq 0}\frac{2\bar{\alpha}_{t}}{(1-\bar{\alpha}_{t})^{2}}$ and $C_{2}=\sup_{t\geq 0}\frac{1+\bar{\alpha}_{t}}{2\sqrt{\bar{\alpha}_{t}}}$. As $10<\bar{\alpha}_{t}<$ 1, it is easy to check the maximum value of $\frac{1+\bar{\alpha}_{t}}{2\sqrt{\bar{\alpha}_{t}}}$, i.e., $C_{2}$ is achieved when $\bar{\alpha}_{t}$ gets close to $1$. This also aligns with the condition where $\sup\frac{2\bar{\alpha}_{t}}{(1-\bar{\alpha}_{t})^{2}}$ is reached. Thus, $\left(1+\frac{1+\bar{\alpha}_{t}}{2\sqrt{\bar{\alpha}_{t}}}\right)M\leq(1+C_{2})M=C_{3}M$ for some constants $C_{3}<2$, since $C_{2}\to 1$ when $\bar{\alpha}_{t}\to 1$. As a result, we have a constant $C>0$ leading to the upper bound as
@@
$|\left\|s_{t_{2}}({\mathbf{x}})\right\|-\left\|s_{t_{1}}({\mathbf{x}})\right\||$ $=\frac{|\left\|s_{t_{2}}({\mathbf{x}})\right\|^{2}-\left\|s_{t_{1}}({\mathbf{x}})\right\|^{2}|}{(\left\|s_{t_{2}}({\mathbf{x}})\right\|+\left\|s_{t_{1}}({\mathbf{x}})\right\|)}$
@@
From Lemma K.4, we have: there exists $C^{\prime}>0$ and $T_{0}>0$ such that for all $t\geq T_{0}$: $||s_{t}({\mathbf{x}})||>C^{\prime}||{\mathbf{x}}||$. Applying this to $t_{1}$, $t_{2}$, and $\xi$ with $\delta>0$, we find that there exists a $C^{\prime}$ such that:
@@
$|\left\|s_{t_{2}}({\mathbf{x}})\right\|-\left\|s_{t_{1}}({\mathbf{x}})\right\||$ $\geq\frac{CK_{2}C^{\prime}\left\|{\mathbf{x}}\right\||t_{2}-t_{1}|}{2C^{\prime}\left\|{\mathbf{x}}\right\|}$
@@
According to Proposition K.5, for two distinct noise levels $t_{1}$ and $t_{2}$, we would obtain different score norms for the same sample ${\mathbf{x}}\in{\mathcal{X}}$. Hence, we consider that different $t^{*}$ would change the score norm of the same sample ${\mathbf{x}}$. In addition, we also imply that the score norms of the same sample ${\mathbf{x}}$ differ if and only if the noise levels $t^{*}$ input to the score network $s_{\theta}$ are different. This is because the score network only takes two input arguments $s_{\theta}({\mathbf{x}},t^{*})$, which is sample ${\mathbf{x}}$ and noise level $t^{*}$. For a natural example ${\mathbf{x}}$, we treat it as a sample with no adversarial perturbation, such that $\epsilon=0$. From Figure 2, cleaner samples typically exhibit smaller score norms, which means the data points are close to the high data density region (Song & Ermon, 2019). Then, we should not diffuse the natural data further with a large $t^{*}$ to retain the original semantic information. Samples subjected to stronger perturbations exhibit larger score norms, which means the adversarial examples are far from the high data density area. We then inject certain amounts of Gaussian noise to the sample to cover its adversarial perturbation during forward diffusion. We identify the score norm as a metric to assess the sample’s proximity to the original data distribution $p({\mathbf{x}})$.

### v2

@@ Appendix I Ablation Study on Bias Term
$p_{t}({\mathbf{x}})=\int_{{\mathcal{X}}}q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0},$ (10)
@@
$p({\mathbf{x}}_{0}|{\mathbf{x}})=\frac{q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})}{\int q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}.$
@@
$\nabla_{{\mathbf{x}}}\log p_{t}({\mathbf{x}})$ $=\nabla_{{\mathbf{x}}}\log\int q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}$ (by Def. K.1) (11)
$=\frac{\nabla_{{\mathbf{x}}}\int q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}{\int q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}$ (chain rule)
$=\frac{\int\nabla_{{\mathbf{x}}}q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}{\int q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}$ (Leibniz integral rule)
$=\frac{\int\frac{\nabla_{{\mathbf{x}}}q({\mathbf{x}}|{\mathbf{x}}_{0})}{q({\mathbf{x}}|{\mathbf{x}}_{0})}q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}{\int q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}$ $\text{(manipulate }q({\mathbf{x}}|{\mathbf{x}}_{0})\text{)}$
$=\frac{\int(\nabla_{{\mathbf{x}}}\log q({\mathbf{x}}|{\mathbf{x}}_{0}))q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}{\int q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}$ (chain rule)
$=\int(\nabla_{{\mathbf{x}}}\log q({\mathbf{x}}|{\mathbf{x}}_{0}))p({\mathbf{x}}_{0}|{\mathbf{x}})d{\mathbf{x}}_{0}$ $\text{(}p({\mathbf{x}}_{0}|{\mathbf{x}})\triangleq\frac{q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})}{\int q({\mathbf{x}}|{\mathbf{x}}_{0})p({\mathbf{x}}_{0})d{\mathbf{x}}_{0}}\text{)}$
@@
Suppose there exists a constant $K>0$ such that for all $t\geq 0$ and all ${\mathbf{x}}_{t}\in{\mathcal{X}}$, the expected norm of the clean data given ${\mathbf{x}}_{t}$ satisfies: $\mathbb{E}_{{\mathbf{x}}_{0}\sim p({\mathbf{x}}_{0}|{\mathbf{x}})}[\left\|{\mathbf{x}}_{0}\right\|]\leq K\left\|{\mathbf{x}}_{t}\right\|$. Then, there exist constants $0<C<1$ and $T_{0}>0$ such that for all $t\geq T_{0}$:
@@
We then turn to look into the asymptotic behavior of $g(t)$. We have $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. Since $0<1-\beta_{s}<1$ for all $s$, and the sequence is decreasing (as $\beta_{s}$ is increasing), it is easy to check that
@@
$1$ $>C.$
Putting them together, we have $\epsilon\in(0,\epsilon_{1})\cup(\epsilon_{2},\infty)$ that makes $\left\|s_{t}({\mathbf{x}})\right\|\geq C\left\|{\mathbf{x}}\right\|$ hold, given a constant $0<C<1$ in relation to $K$ and $\epsilon$. Setting $T_{0}=T_{\epsilon}$ completes the proof. ∎
@@
Next, under Cauchy-Schwarz inequality: $|\left<{\mathbf{a}},{\mathbf{b}}\right>|\leq\left\|{\mathbf{a}}\right\|\left\|{\mathbf{b}}\right\|$, we have
@@
Let $C_{1}=\sup_{t\geq 0}\frac{2\bar{\alpha}_{t}}{(1-\bar{\alpha}_{t})^{2}}$ and $C_{2}=\sup_{t\geq 0}\frac{1+\bar{\alpha}_{t}}{2\sqrt{\bar{\alpha}_{t}}}$. As $0<\bar{\alpha}_{t}<1$, it is easy to check the maximum value of $\frac{1+\bar{\alpha}_{t}}{2\sqrt{\bar{\alpha}_{t}}}$, i.e., $C_{2}$ is achieved when $\bar{\alpha}_{t}$ gets close to $1$. This also aligns with the condition where $\sup\frac{2\bar{\alpha}_{t}}{(1-\bar{\alpha}_{t})^{2}}$ is reached. Thus, $\left(1+\frac{1+\bar{\alpha}_{t}}{2\sqrt{\bar{\alpha}_{t}}}\right)M\leq(1+C_{2})M=C_{3}M$ for some constants $C_{3}<2$, since $C_{2}\to 1$ when $\bar{\alpha}_{t}\to 1$. As a result, we have a constant $C>0$ leading to the upper bound as
@@
[…]
@@
[…]
@@
[…]
@@
According to Proposition K.5, for two distinct noise levels $t_{1}$ and $t_{2}$, we would obtain different score norms for the same sample ${\mathbf{x}}\in{\mathcal{X}}$. Hence, we consider that different $t^{*}$ would change the score norm of the same sample $\mathbf{x}$}. In addition, we also imply that the score norms of the same sample $\mathbf{x}$} differ if and only if the noise levels $t^{*}$ input to the score network $s_{\theta}$ are different. This is because the score network only takes two input arguments $s_{\theta}({\mathbf{x}},t^{*})$, which is sample $\mathbf{x}$} and noise level $t^{*}$. For a natural example $\mathbf{x}$}, we treat it as a sample with no adversarial perturbation, such that $\epsilon=0$. From Figure 2, cleaner samples typically exhibit smaller score norms, which means the data points are close to the high data density region (Song and Ermon, 2019). Then, we should not diffuse the natural data further with a large $t^{*}$ to retain the original semantic information. Samples subjected to stronger perturbations exhibit larger score norms, which means the adversarial examples are far from the high data density area. We then inject certain amounts of Gaussian noise to the sample to cover its adversarial perturbation during forward diffusion. We identify the score norm as a metric to assess the sample’s proximity to the original data distribution $p({\mathbf{x}})$.